# Constrained least squares simplicial-simplicial regression

### v2

@@ 4.4 Computational cost
To evaluate the computational cost the command benchmark(), from the R package Rfast2 (Papadakis et al., 2023), was used, based on 20 repetitions. Simplicial response values with $D_{r}=(3,5,7,10)$ and simplicial predictor values with $D_{p}=3$ were generated from a Dirichlet distribution and the sample sizes considered were $n=(500,1000,5000,10000,20000)$. The speedup factors of the SCLS compared to TFLR are presented in Table 4. The ratios clearly depict that the time required to fit the TFLR model is significantly higher compared to that of the SCLS model.
@@
codalm implementation Self implementation
@@ 5.4 Predictive performance of the power transformed SCLS model
The FADN and the education data sets were further used to test the performance of the SCLS model when applied to the power transformed data using Eq. (5) employing the 20 times repeated 10-fold CV protocol. Figure 4 presents the average predictive KLD measures as a function of the $\alpha$-values. Evidently, this strategy was not proved prosperous for the FADN data, but it was beneficiary in the case of the Education data set.
@@ 6 Conclusions
Closing this paper we would like to pose some possible research directions and suggestions. a) Exploitation of the relationship between compositional and directional data. By taking the square root both variables transform into directional data for which spherical regression models exist and in order to ensure that the fitted values will lie within the simplex the folded Kent model (Scealy and Welsh, 2014) may be employed. b) Change of the loss function to the $L_{1}$ norm, i.e. constrained minimization of the sum of the absolute errors. c) Investigation of the ensemble learning in the $\alpha$-SCLS model. Instead of selecting one value of $\alpha$ combine the fitted values of many models produced by different values of $\alpha$. d) The possibility to add non-linear effects in the model, regardless of the loss function used, to increase the flexibility of the model and escape the assumption of linear relationship.
@@
Instead of calling the function ols.compcomp() from the R package Compositional multiple times, a function was created minimizing the necessary to perform computations. The vector ${\bf b}_{0}$ is always the same, and so is the matrix $\bf D$ since even by permuting the rows of the simplicial predictor $\bf X$ the cross-product ${\bf X}^{\top}{\bf X}$ remains the same. The only thing that changes is the $\bf d$ vector. Secondly, Eq. (13) was computed excluding the trace of the matrix ${\bf Y}{\bf Y}^{\top}$ as this is constant.

### v3

@@ 4.4 Computational cost
To evaluate the computational cost the command benchmark(), from the R package Rfast2 (Papadakis et al., 2023), was used, based on 20 repetitions. Simplicial response values with $D_{r}=(3,5,7,10)$ and simplicial predictor values with $D_{p}=3$ were generated from a Dirichlet distribution and large sample sizes were considered, $n=(500,1000,5000,10000,20000)$. The speedup factors of the SCLS compared to TFLR are presented in Table 4. The ratios clearly depict that the time required to fit the TFLR model is significantly higher compared to that of the SCLS model.
@@
codalm implementation Self implementation of EM algorithm
@@ 5.4 Predictive performance of the power transformed SCLS model
The FADN and the education data sets were further used to test the performance of the SCLS model when applied to the power transformed data using Eq. (5) employing the 10-fold CV protocol repeated 20 times. Figure 4 presents the average predictive KLD measures as a function of the $\alpha$-values. Evidently, this strategy was not proved prosperous for the FADN data, but it was beneficiary in the case of the Education data set.
@@ 6 Conclusions
Closing this paper we would like to pose some possible research directions and suggestions. We live in the era of big data and algorithms ought to be customizable to treat such cases. The SCLS implementation in the R package Compositional addresses this issue by accepting Filebacked Big Matrices. Another option would be to ”cut” the data in chunks and via parallel computing perform the SCLS model in each chunk and then combine the results in a meta-analytic way. An alternative, and promising, strategy would be to perform subdata selection (Chasiotis and Karlis, 2024).
The SURE idea may be used for other multivariate regression models or loss functions as well. For instance, extending the $L_{1}$ and $L_{2}$ norm regression models to the general $L_{p}$ regression model (Money et al., 1982). Other options include the substitution of the the KL divergence by the Jensen-Shannon divergence or any other $\phi$-divergence. The common ground in these aforementioned extensions is the set of constraints imposed on the matrix of regression coefficients $\bf B$ and the linear link between the response and predictor variables.
Finally some more suggestions are the following. (a) Investigation of any possible benefit of the power transformation (5) to the simplicial predictor(s). This could yield more more accurate predictions, but at the cost of increased computational cost. b) Exploitation of the relationship between compositional and directional data. By taking the square root both variables transform into directional data for which spherical regression models exist and in order to ensure that the fitted values will lie within the simplex the folded Kent model (Scealy and Welsh, 2014) may be employed. c) Investigation of the ensemble learning in the $\alpha$-SCLS model. Instead of selecting one value of $\alpha$ combine the fitted values of many models produced by different values of $\alpha$. d) The possibility to add non-linear effects in the model, regardless of the loss function used, to increase the flexibility of the model and escape the assumption of linear relationship. For instance, the use of neural networks, or local polynomial regression with simplicial constraints imposed on the coefficients.
@@
Instead of calling the function scls() from the R package Compositional multiple times, a function was created minimizing the necessity to perform computations. The vector ${\bf b}_{0}$ is always the same, and so is the matrix $\bf D$ since even by permuting the rows of the simplicial predictor $\bf X$ the cross-product ${\bf X}^{\top}{\bf X}$ remains the same. The only thing that changes is the $\bf d$ vector. Secondly, Eq. (13) was computed excluding the trace of the matrix ${\bf Y}{\bf Y}^{\top}$ as this is constant.